Solve the following algebraically. Please show the work
t^(2/3)= 4
√(5&x) + 1 = 3
2/3= 2- (5x-3)/(x-1)
√(t+12) - t=0
The volume of a cube is given V= s^3, where s is the length of a side. Find the length of a side of a cube if the volume is 700 〖in〗^3. Round the answer to three decimal places

Answer by checkley77(12844)

t^(2/3)=4
cubert(t^2)=4
now cube each side
t^2=4^3
t^2=64
t=sqrt64
t=8 answer.
proof
t^2/3=4
8^2/3=4
cubert8^2=4
cubert64=4
4=4